# Sharp Time Decay and Regularization

### v1

@@ 1.1 Main theorems
with respect to the convolution, i.e., $k*\ell=1$. It is worth mentioning that there are different types pf kernels that define other types of fractional derivative, and it would be interesting to analyze, see [1, 43, 51].
The operator $\mathcal{L}$ is a linear diffusion operator of local or nonlocal type. We refer to Section 1.2 for a precise description, and to Section 1.3 for some examples. The solutions that we consider here are defined in the Hilbert space $H^{*}(\Omega)$, which is the topological dual space of $H(\Omega)$, the domain of the quadratic form associated to the operator $\mathcal{L}$:
@@
($H^{*}$-solutions). Let $u\in L^{2}((0,T):H^{*}(\Omega))$, then $u$ is a $H^{*}$-solution of (CPME) if the following hold:
@@
Our first main result is a comparison principle for $H^{*}$-solutions, which has its own interest, and it is also a key ingredient for proving the rest of the main results presented in this paper. To the best of our knowledge, this comparison result is new and a complete proof can be found in Section 3.2, based on the discretized problem studied in Section 2. For the sake of completeness, together with the comparison principle, we state existence and uniqueness result proven by Akagi in [1] and by Li and Liu in [37].
@@
The proof of the comparison result is based on a suitable discretization in the spirit of [37], together with a clever limiting process: we prove comparison first for the Euler implicit scheme associated to the $H^{*}$ “gradient flow structure” of (CPME). We study the discretized problem in Section 2, where we show existence, uniqueness of solutions and a T-contractivity property in a weighted $L^{1}$ space, cf. Theorem 2.4. The T-contractivity estimate at the discretized level, allows us to deduce a discrete comparison principle, cf. Corollary 2.5. This last result goes all the way through the limiting process and allows to prove the desired comparison principle in the continuous framework, cf. Section 3.
A direct – and important – consequence of the comparison principle and of the time-scaling of the equation are the Benilan-Crandall time monotonicity formulae that we present below. Let us emphasize that achieving any time monotonicity in PDEs involving the Caputo time derivative is considerably more challenging than with the classical time derivative, due to the “memory effect”: when $D^{\alpha}_{t}f(t)\leq 0$, we can only deduce a stability estimate, $f(t)\leq f(0)$, rather than the “full” monotonicity, i.e. $f(t)\leq f(t_{0})$ for every $t_{0}<t$.
@@
Section 4 is devoted to the proof of the new $L^{p}-L^{\infty}$ smoothing effects and boundary estimates. As already mentioned, the only $L^{p}-L^{\infty}$ regularizing effects involving the Caputo time derivative present in the literature have been established for the Heat equation in the whole space $\mathbb{R}^{N}$, see [34, Theorem 3.3.]: the proof relies on the representation formula and on the fundamental solution which is unbounded and has an explicit form. In our setting the proof is much more involved, due to the lack of a representation formula and the nonlinearity.
@@
Surprisingly, the decay rate $t^{-\frac{\alpha}{m}}$ of the $L^{\infty}$-norm for $\alpha\in(0,1)$ presented in (6) is significantly different than in the case $\alpha=1$. For the porous medium equation with classical time derivative, the existence of the Friendly Giant provides that the decay of the $L^{\infty}$-norm matches with the time scaling:
@@
with a universal constant $K>0$, see [15, 17]. In addition, the smoothing effects for small times provided in [13, 17] for the porous medium reads as
$\|u(t)\|_{\infty}\leq\overline{\kappa}\,\|u_{0}\|_{p}^{2sp\vartheta_{p}}t^{-N\vartheta_{p}}\qquad\forall t>0,\quad\alpha=1,\quad m>0\,,$ (8)
where $u_{0}\in L^{p}(\Omega)$ with $p>\frac{N(1-m)}{2s}$ and $\vartheta_{p}=(2sp-N(1-m))^{-1}$. However, for $\alpha\in(0,1)$, the time scaling of (CPME) would suggest a $L^{\infty}$ decay of the form $t^{-\frac{\alpha}{m-1}}$, but this is not the case, as (6) shows.
@@ 1.2 Assumptions and functional setting of the operator $\mathcal{L}$
for some $s\in(0,1]$. Assumption (G1) guarantees the compactness of the operator $\mathcal{L}^{-1}$ in $L^{2}(\Omega)$, see [11, Proposition 5.1]. As a consequence, the operator $\mathcal{L}^{-1}$ has a discrete spectrum and a $L^{2}(\Omega)$ orthonormal basis of eigenfunctions that we denote by $\left(\lambda_{k}^{-1},\Phi_{k}\right)_{k\geq 1}$. Accordingly, we can express
@@ 1.3 Examples of $\mathcal{L}$
The assumptions that we state above describe a wide class of operators of local and nonlocal type. Let us enumerate some of the most important ones here. For a more extended list of examples we refer to [13] and [18, Section 3].
@@
Other examples: As explained in [18, Section 3], our theory may also be applied to: i) the operators above with uniformly elliptic, $C^{1}$ and symmetric coefficients; ii) sums of two fractional operators; iii) the sum of the classical Laplacian and a nonlocal operator kernel; iv) Schrödinger equations for non-symmetric diffusions; v) gradient perturbation of restricted fractional Laplacians.
@@
Recall that $H_{0}^{s}(\Omega)=H^{s}(\Omega)$ when $s\leq 1/2$, and hence the dual coincide. If $s=1/2$, then $H^{s}_{00}(\Omega)\subsetneq H^{s}_{0}(\Omega)$, which implies $H^{-s}(\Omega)\subsetneq\left(H^{s}_{00}(\Omega)\right)^{*}$. See [15, Sec. 7.7] for more details. This identification holds for any operator $\mathcal{L}$ with discrete spectrum such that $0<\lambda_{k+1}/\lambda_{k}\leq\Lambda_{0}$ for some $\Lambda_{0}>0$.
@@ 2 Discretized problem
In this section we will provide some properties of the discretization of (CPME) and then deduce the comparison principle. Several discretizations of the Caputo derivative have been proposed through the years, see for instance [27, 37, 38]. We choose the one in [37] based on a deconvolution scheme, since it is enough to obtain the convergence results needed in this work.
@@
[37, Section 3.] Let $\tau=T/N>0$ be the time-step and $t_{k}=k\tau$ with $0\leq k\leq N$ points of the interval $(0,T)$. Let us consider the Euler implicit scheme for (CPME):
@@ 2.1 Existence and uniqueness for the discretized problem
The idea of the proof is the same of the nonlinear theory of Brezis for gradient flows [20], see also Komura [36] and the excellent lecture notes [4]. Due to the convexity of ${\mathcal{E}}$ we can construct recursively the next function $U_{k+1}$ with $k\geq 0$ by the minimizing movements scheme, namely, for any $k\geq 0$
@@ 2.2 Discrete Comparison Principle
see [11, Theorem 7.1. and 7.2.] for more details.
Let us remark that for nonnegative functions $u$ the belong to $H^{*}(\Omega)$ also belongs to $L^{1}_{\Phi_{1}}(\Omega)$:
@@
The theorem below follows the formal proof of the classical T-contractivity estimate in $L^{1}_{\Phi_{1}}$ [13, Theorem 2.4.] and the rigorous techniques used to show the $L^{1}$-contractivity for the linear equation with Caputo derivative [51, Theorem 7.1.]. Both techniques require extra regularity on the time derivative, more precisely $\partial_{t}u\in L^{1}((0,T):L^{1}_{\Phi_{1}}(\Omega))$ or $D^{\alpha}_{t}u\in L^{\infty}((0,T):L^{1}(\Omega))$, respectively. Since we only expect, a-priori, that $D^{\alpha}_{t}u\in L^{2}((0,T):H^{*}(\Omega))$, we first prove T-contractivity at the discrete level, and this allows to show comparison principle as $\tau\to 0$.
@@
(Discrete T-contractivity in $L^{1}_{\Phi_{1}}(\Omega)$). Let $m>0$ and assume (G1) and (K). Consider $\{U_{k}\}_{k\geq 0},\{V_{k}\}_{k\geq 0}\subset L^{1+m}(\Omega)\cap H^{*}(\Omega)$ two sequence of solutions for (17). Then,

### v3

@@ 1.1 Main theorems
with respect to the convolution, i.e., $k*\ell=1$. It is worth mentioning that there are different types of kernels that define distinct fractional derivatives, and it would be interesting to analyze, see [1, 67, 76].
The operator $\mathcal{L}$ is a linear diffusion operator of local or nonlocal type. We refer to Section 1.3 and 1.4 for a more precise description and examples. We briefly recall here the two main assumptions: the inverse operator $\mathcal{L}^{-1}:L^{2}(\Omega)\rightarrow L^{2}(\Omega)$ is represented via a kernel $\mathbb{G}_{\Omega}$:
$\mathcal{L}^{-1}u(x):=\int_{\Omega}\mathbb{G}_{\Omega}(x,y)u(y)\,{\rm d}y\,.$
The Green function $\mathbb{G}_{\Omega}$ satisfies
$0\leq\mathbb{G}_{\Omega}(x,y)\leq\frac{c_{1,\Omega}}{|x-y|^{N-2s}}\,,$ (G1)
for some $s\in(0,1]$. Assumption (G1) guarantees the compactness of the operator $\mathcal{L}^{-1}$ in $L^{2}(\Omega)$, see [13, Proposition 5.1]. As a consequence, the operator $\mathcal{L}^{-1}$ has a discrete spectrum and a $L^{2}(\Omega)$ orthonormal basis of eigenfunctions that we denote by $\left(\lambda_{k}^{-1},\Phi_{k}\right)_{k\geq 1}$. We moreover assume that $\mathcal{L}$ satisfies the following weak form of Kato’s inequality:
$\int_{\Omega}\mathcal{L}\left[h(u)\right]\Phi_{1}\,{\rm d}x\leq\int_{\Omega}h^{\prime}(u)\mathcal{L}u\,\Phi_{1}\,{\rm d}x\,,$ (K)
for any convex function $h\in C^{1}(\mathbb{R})$ with $h(0)=0$, any function $u\in L^{1}_{\rm loc}(\Omega)$, and $\Phi_{1}\geq 0$ the first eigenfunction of $\mathcal{L}$.
The solutions that we consider here are defined in the Hilbert space $H^{*}(\Omega)$, which is the topological dual space of $H(\Omega)$, the domain of the quadratic form associated to the operator $\mathcal{L}$:
@@
($H^{*}$-solutions). Let $u\in L^{2}((0,T):H^{*}(\Omega))$, then $u$ is a $H^{*}$-solution of (CPME) if the following holds:
@@
Our first main result is a comparison principle for $H^{*}$-solutions, which has its own interest, and it is also a key ingredient for proving the rest of the results presented in this paper. To the best of our knowledge, this comparison result is new and a complete proof can be found in Section 3.2, based on the discretized problem studied in Section 2. For the sake of completeness, together with the comparison principle, we state existence and uniqueness result proven by Akagi in [1] and by Li and Liu in [54].
@@
The proof of the comparison result is based on a suitable discretization in the spirit of [54], together with a clever limiting process: we first prove comparison for the Euler implicit scheme associated to the $H^{*}$ “gradient flow structure” of (CPME). We study the discretized problem in Section 2, where we show existence, uniqueness of solution, and a T-contractivity property in a weighted $L^{1}$ space, cf. Theorem 2.4. The T-contractivity estimate at the discretized level, allows us to prove a discrete comparison principle, cf. Corollary 2.5. This last result goes all the way through the limiting process and allows to prove the desired comparison principle in the continuous framework, cf. Section 3. This method diverges from conventional PDE techniques. Typically, the limit solution from a discretization process is obtained through compactness arguments, and loses the well-ordering property. In our case, however, the uniqueness of gradient flow solution allows to keep the desired comparison. Despite the fact that classical solutions and viscosity solutions are the natural classes of solutions where the comparison principle makes sense pointwise [56], we prove the result above for solutions in the weak space $H^{*}(\Omega)\cap L^{1+m}(\Omega)$, which introduces an additional level of difficulty.
A direct – and important – consequence of the comparison principle and of the time-scaling are the Benilan-Crandall time monotonicity formulae that we present below. Let us emphasize that achieving any time monotonicity in PDEs involving the Caputo time derivative is considerably more challenging than with the classical time derivative, due to the “memory effect”: when $D^{\alpha}_{t}f(t)\leq 0$, we can only deduce a stability estimate, $f(t)\leq f(0)$, rather than the “full” monotonicity, i.e. $f(t)\leq f(t_{0})$ for every $t_{0}<t$.
@@
Section 4 is devoted to the proof of the new $L^{p}-L^{\infty}$ smoothing effects and boundary estimates. As already mentioned, the only $L^{p}-L^{\infty}$ regularizing effects involving the Caputo time derivative present in the literature have been established for the Heat equation. We refer to [50, Theorem 3.3.] in the case of unbounded domains: there, the proof relies on the representation formula and the explicit fundamental solution. For the Heat equation on bounded domains, smoothing effects for the time fractional problem can be inherited from the classical problem with $\alpha=1$, combining the estimates in [32, Theorem 6.5] with [46, Proposition 2.2.2]. In our setting the proof is much more involved, due to the nonlinearity and the lack of a representation formula.
@@
Surprisingly, the decay rate $t^{-\frac{\alpha}{m}}$ in (6) is significantly different than in the case $\alpha=1$. For the porous medium equation with classical time derivative, the existence of the Friendly Giant - separate-variable solution of the form $S(x)\,t^{-\frac{1}{m-1}}$ provides that the decay of the $L^{\infty}$-norm matches with the time scaling:
@@
with a universal constant $K>0$, see [18, 20]. In addition, the smoothing effects for small times provided in [16, 20] for the classical porous medium reads as
$\|u(t)\|_{\infty}\leq\overline{\kappa}\,\|u_{0}\|_{p}^{2sp\vartheta_{p}}t^{-N\vartheta_{p}}\qquad\forall 0<t<1\normalcolor,\quad\alpha=1,\quad m>0\,,$ (8)
where $u_{0}\in L^{p}(\Omega)$ with $p>\frac{N(1-m)}{2s}$ and $\vartheta_{p}=(2sp-N(1-m))^{-1}$. Interestingly, the decay (6) corresponds to the decay of solutions to the equation with $\alpha=1$ and $\Omega=\mathbb{R}^{N}$, see [15, Theorem 3.1].
@@
Remark 1 (The case of unbounded domains).
A closer inspection of the proofs shows that the results proven on Theorems 1.2, 1.3 and 1.4 can be generalized to unbounded domains including +$\mathbb{R}^{N}_{+}$ and $\mathbb{R}^{N}$ with minimal effort. The existence and uniqueness theory of fractional gradient flows, the comparison principle and the time monotonicity carry over essentially unchanged. The smoothing effects follows by decomposing the unbounded domain into a ball of a specific radius and its complement. In this way, it is possible to obtain the same $L^{p}-L^{\infty}$ estimate in unbounded domains, following the method in [15]. Alternatively, following [42], one can use a sequence of solutions to (CPME) on expanding balls, to which all theorems here apply and provide uniform estimates respect to the size of the ball, and pass to the limit.
1.2 Physical motivation of the model
The classical Porous Medium Equation with $\alpha=1$ and $m>1$ is named after its usefulness to describe the behaviour of an ideal gas in a homogeneous porous medium according to the works of Leibenzon [53] and Muskat [61]. Similarly, this model also provides a good description of a fluid, typically water, penetrating in a porous medium, such as soil or brick. However, this equation has been used widely to model different types of nonlinear diffusion processes such as nonlinear heat transfer and population dynamics, see [70] and the references therein. We recall from [63, 70] the derivation of the physical model for the groundwater flow in unsaturated porous media.
At a macroscopic level, let us consider $u(t,x)$ to be the fluid concentration at point $x$ at time $t$, and let the vector $\textbf{q}(t,x)$ represent the flux, which denotes the amount and direction of fluid across the unit surface in a unit time. According to the mass conservation property, the function $u$ satisfies the continuity equation
$\partial_{t}u+\nabla\cdot\textbf{q}=0\,.$ (12)
In 1856, Darcy [34] showed empirically that the flux $q$ is proportional to the gradient of the pressure
$\textbf{q}=-\frac{\rho k}{\mu}\nabla p\,,$ (13)
where $\rho>0$ is the density of the fluid, $\mu>0$ is the viscosity, $k>0$ is the permeability and $p$ is the pressure function. Let us remark that, Darcy’s Law can also be derived rigorously by averaging Navier-Stokes equations for fluids in a porous medium [75]. In most cases and with most fluids, the pressure $p$ can be expressed as a monotone function of the fluid concentration $u$, such that $p=p(u)$ and hence
$\textbf{q}=-\frac{\rho k}{\mu}\frac{{\rm d}p}{{\rm d}u}\nabla u=:-D(u)\nabla u\,,$ (14)
where $D(u)$ is known as the diffusivity. Implementing this definition in the continuity equation lead us to the following equation
$\partial_{t}u=\nabla\cdot\left(D(u)\nabla u\right)\,,$ (15)
known in the literature as Richards equation [65] and Boussinesq’s equation [23]. The selection of the pressure $p(u)$ is frequently chosen to fit the experimental results. In particular, one of the most common is the Brooks-Correy function [25]
$p(u)\sim u^{m}\,,$ (16)
where the value $m$ varies depending on the soil type under examination and the distribution of the pores within it.
Let us mention that the Fast Diffusion Equation, i.e. $m\in(0,1)$, arises in the gas-kinetics model of Barenblatt-Zel’dovich-Kompaneets [8], and also in plasma physics where high temperatures alter standard physical laws (Fourier’s Law). In this context, the Okuda-Dawson scaling [62] leads to a singular diffusivity $D(u)\sim u^{-1/2}$ in (15). See the pioneering work of Berryman and Holland [9].
Recent studies have observed that these models are not accurate enough when addressing certain porous media, such as building materials [44] and zeolite [37]. In this cases, fluid particles can be trapped in some regions for a significant amount of time delaying the transport of the moisture. Following [63], let us show a deterministic derivation of the Caputo Porous Medium Equation for this type of phenomena.
Given a fluid in a porous medium which flow behaves according to (12), let us assume that certain physical phenomena cause fluid particles to become trapped in specific regions for a duration denoted by $\tau$. As a result, only a part of the instantaneous flux at time $t$ should contribute to the change in the concentration. But we should also consider the released particles that were trapped in the previous time $t-\tau$. Consequently, we can rewrite (12) as delayed equation with corresponding weights, denoted by $w_{0}$ and $w_{1}$,
$\partial_{t}u(t)=-\left(w_{0}\nabla\cdot\textbf{q}(t)+w_{1}\nabla\cdot\textbf{q}(t-\tau)\right)\,.$
Following this idea, we can build the model for the case when particles are trapped in different periods of time $0\leq\tau_{1}\leq\dots\leq\tau_{n}<t$ with their associated weights $w_{1},\dots,w_{n}$
$\partial_{t}u(t)=-\sum_{i=1}^{n}w_{i}\nabla\cdot\textbf{q}(t-\tau_{i})\,.$
By incorporating the weight density $w=w(\tau)$, we can generalize the trapping phenomena to a continuous distribution of periods. Therefore, since $w_{i}=w(\tau_{i})(\tau_{i}-\tau_{i-1})$ we obtain
$\partial_{t}u(t)=-\sum_{i=1}^{n}w(\tau_{i})\nabla\cdot\textbf{q}(t-\tau_{i})(\tau_{i}-\tau_{i-1})\quad\longrightarrow\quad\partial_{t}u(t)=-\int_{0}^{t}w(\tau)\nabla\cdot\textbf{q}(t-\tau)\,{\rm d}\tau\,,$
which lead us to the non-local integro-differential equation by a change of variable:
$\partial_{t}u(t)=-\int_{0}^{t}w(t-\tau)\nabla\cdot\textbf{q}(\tau)\,{\rm d}\tau\,.$ (17)
We consider now the family of weight densities $w_{\alpha}$ with $\alpha\in(0,1)$, satisfying $w_{\alpha}\rightarrow\delta_{0}$ in the sense of distributions when $\alpha\rightarrow 1$. Note that when $\alpha\rightarrow 1$ the particles are being trapped for shorter time and the memory-effect is decreasing until equation (12) is recovered. As we consider $w_{\alpha}$ to be a distribution, let us modified the kernel in (17) to deal with functions. Using the approximation of Dirac’s delta given by
$w_{1}=\delta_{0}=\lim\limits_{h\rightarrow 0}\frac{\chi_{[0,h]}}{h}$
with the limit in the weak sense, then
$\nabla\textbf{q}(t)=\int_{0}^{t}w_{1}(t-\tau)\nabla\cdot\textbf{q}(\tau)\,{\rm d}\tau=\lim\limits_{h\rightarrow 0}\int_{t-h}^{t}\nabla\cdot\textbf{q}(\tau)\,{\rm d}\tau=\frac{\rm d}{\,{\rm d}t}\int_{0}^{t}\nabla\cdot\textbf{q}(\tau)\,{\rm d}\tau\,.$
Hence, we rewrite equation (17) in terms of a function $k_{\alpha}$ with $k_{1}\equiv 1$ as follows
$\partial_{t}u(t)=-\frac{\rm d}{\,{\rm d}t}\int_{0}^{t}k_{\alpha}(t-\tau)\nabla\cdot\textbf{q}(\tau)\,{\rm d}\tau\,.$ (18)
It this framework, there are plenty of kernels $k_{\alpha}$ to be consider, see [1, 67, 76]. But, the simplest one is to consider a power function with a normalizing constant $k_{\alpha}(\tau)=\Gamma(\alpha)^{-1}\tau^{\alpha-1}$,
$\partial_{t}u(t)=-\frac{1}{\Gamma(\alpha)}\frac{\rm d}{\,{\rm d}t}\int_{0}^{t}(t-\tau)^{\alpha-1}\nabla\cdot\textbf{q}(\tau)\,{\rm d}\tau=:-\prescript{}{RL}{D_{t}^{1-\alpha}}\big{(}\nabla\cdot\textbf{q}(t)\big{)}\,.$ (19)
Note that we have obtained the Riemann-Liouville fractional derivative in the right hand side. Given suitable regularity conditions, see [43, Theorem 2.22], we can apply Riemann-Liouville integral in both sides of (19) to obtain the following equation involving the Caputo derivative:
$D_{t}^{\alpha}u(t)=-\nabla\cdot\textbf{q}(t)\,.$ (20)
This equation corresponds to (CPME) with flux from Darcy’s Law (14) and pressure (16), specifically when $s=1$.
@@ 1.3 Assumptions and functional setting of the operator $\mathcal{L}$
for some $s\in(0,1]$. Assumption (G1) guarantees the compactness of the operator $\mathcal{L}^{-1}$ in $L^{2}(\Omega)$, see [13, Proposition 5.1]. As a consequence, the operator $\mathcal{L}^{-1}$ has a discrete spectrum and a $L^{2}(\Omega)$ orthonormal basis of eigenfunctions that we denote by $\left(\lambda_{k}^{-1},\Phi_{k}\right)_{k\geq 1}$. Accordingly, we can express
@@ 1.4 Examples of $\mathcal{L}$
The assumptions that we state above describe a wide class of operators of local and nonlocal type. Let us enumerate some of the most important ones here. For a more extended list of examples we refer to [16] and [21, Section 3].
@@
Other examples: As explained in [21, Section 3], our theory may also be applied to: i) the operators above with uniformly elliptic, $C^{1}$ and symmetric coefficients; ii) sums of two fractional operators; iii) the sum of the classical Laplacian and a nonlocal operator kernel; iv) Schrödinger equations for non-symmetric diffusions; v) gradient perturbation of restricted fractional Laplacians.
@@
Recall that $H_{0}^{s}(\Omega)=H^{s}(\Omega)$ when $s\leq 1/2$, and hence the dual coincide. If $s=1/2$, then $H^{s}_{00}(\Omega)\subsetneq H^{s}_{0}(\Omega)$, which implies $H^{-s}(\Omega)\subsetneq\left(H^{s}_{00}(\Omega)\right)^{*}$. See [18, Sec. 7.7] for more details. This identification holds for any operator $\mathcal{L}$ with discrete spectrum such that $0<\lambda_{k+1}/\lambda_{k}\leq\Lambda_{0}$ for some $\Lambda_{0}>0$.
@@ 2 Discretized problem
In this section we will provide some properties of the discretization of (CPME) and then deduce the comparison principle. Several discretizations of the Caputo derivative have been proposed through the years, see for instance [36, 54, 55]. The discretization that works best for our convergence analysis is the one proposed in [54], which is based on a deconvolution scheme.
@@
[54, Section 3.] Let $\tau=T/N>0$ be the time-step and $t_{k}=k\tau$ with $0\leq k\leq N$ points of the interval $(0,T)$. Let us consider the Euler implicit scheme for (CPME):
@@ 2.1 Existence and uniqueness for the discretized problem
The idea of the proof is the same of the nonlinear theory of Brezis for gradient flows [24], see also Komura [52], the celebrated JKO scheme [49] and the excellent lecture notes [4]. Due to the convexity of ${\mathcal{E}}$ we can construct recursively the next function $U_{k+1}$ with $k\geq 0$ by the minimizing movements scheme, namely, for any $k\geq 0$
@@ 2.2 Discrete Comparison Principle
see [13, Theorem 7.1. and 7.2.] for more details.
Let us remark that nonnegative functions $u$ belonging to $H^{*}(\Omega)$ also belong to $L^{1}_{\Phi_{1}}(\Omega)$
@@
The next theorem follows the formal proof of the classical T-contractivity estimate in $L^{1}_{\Phi_{1}}$ [16, Theorem 2.4.] and the rigorous techniques used to show the $L^{1}$-contractivity for the linear equation with Caputo derivative [76, Theorem 7.1.]. Both techniques require extra regularity on the time derivative, more precisely $\partial_{t}u\in L^{1}((0,T):L^{1}_{\Phi_{1}}(\Omega))$ or $D^{\alpha}_{t}u\in L^{\infty}((0,T):L^{1}(\Omega))$, respectively. Since we only expect, a-priori, that $D^{\alpha}_{t}u\in L^{2}((0,T):H^{*}(\Omega))$, we first prove T-contractivity at the discrete level, and then comparison principle in the limit $\tau\to 0$.
@@
(Discrete T-contractivity in $L^{1}_{\Phi_{1}}(\Omega)$). Let $m>0$ and assume (G1) and (K). Consider $\{U_{k}\}_{k\geq 0},\{V_{k}\}_{k\geq 0}\subset L^{1+m}(\Omega)\cap H^{*}(\Omega)$ two sequences of solutions for (26). Then,